OpenStudy (anonymous):

1st half year sales volume 5 oz is 20% 10oz is 30% 15oz is 50% 2nd half year sales volume: 5 oz is 30% 10oz is 40% 15oz is 30% What is average oz sold for that year?

OpenStudy (anonymous):

Alright, so the average of a quantity is the summation of each individual quantity multiplied by it's probability. In this case, we are dealing with an interesting set up, so we'll take it apart a bit. What we'll do is take each quantity (for example, 5) and multiply it by the percent next to it. Do this for each term. Sum them up. Then divide by the total percents. This may be a little confusing, so I'll demonstrate if you want.

OpenStudy (anonymous):

This is how i calculated mine: 5 oz (20) + 5oz (30)= 250 10 oz (30)+ 10 (40)= 700 15 oz (50) + 15 (30)= 1200 Total is 2150/6= 358.33?

OpenStudy (anonymous):

Very close. Your first step was correct. You need to divide by the summation of the weighting factors (20+30+50+30+40+30 = 200)

OpenStudy (anonymous):

ohh ok so 2150/200?

OpenStudy (anonymous):

Yeah. Always ask if your answer makes sense. You are averaging between three values, 5, 10, and 15. Your answer will be somewhere in the range between 5 and 15.

OpenStudy (anonymous):

ya i knew that i was almost there but not quite..but in this case i take the total weighted values for each year and divided by the total for both years to get my average for those years right

OpenStudy (anonymous):

Yeah. You're given the percents of each sold in two different half years, which is why the set up is a little odd. So, each half year has a maximum percent of 100, making 200% the maximum for a full year.

OpenStudy (anonymous):

For this question...can I show you how I would do it?

OpenStudy (anonymous):

A new company will release 3 new products (A, B, C). 20% of buyers will buy A, 55% will buy B and 25% will buy C. People that buy A will buy 1.5 units, B 2.5 and C .75. What percent of total sales is Product B?

OpenStudy (anonymous):

PA 20 (1.5)= 30 PB 55 (2.5)= 137.5 PC 25 (.75)= 18.75 Total= 186.25 therefore, 137.5/186.25 is approx 74%

OpenStudy (anonymous):

Looks fine to me!

OpenStudy (anonymous):

Great- thank you for verifying...it was very helpful

OpenStudy (anonymous):

What about this? I have an idea how to get it but not sure how to properly show it: Out of a random sample of 500, 75 like oranges. How much more sample should be taken in order to reach the no 100 who like oranges

OpenStudy (anonymous):

So, there isn't REALLY a solid way to answer this question, as we have no information to base where our current results are distributed statistically over the population. However, if we assume that our 500 results follow the normal population distribution, it's a simple ratio. 75:500 :: 100:X

OpenStudy (anonymous):

So, 500/75= 15% like orange To find 100 people in total that like oranges I need to do 100/.15= 666.67 so 666.67-500= 166.67 more people?

OpenStudy (anonymous):

That'll do it as well :)

OpenStudy (anonymous):

The one caution I'd use is that you can't question 0.67 of a person, so you'll likely want to round to a whole person.

OpenStudy (anonymous):

how would you calculate this question?

OpenStudy (anonymous):

You did it correctly, it just that you can either ask a person or not, you can't ask 2/3 of a person. Therefore, a decimal doesn't really make sense. You may want to check with your teacher about rounding, however. They may want the decimal, and it's really up to them.

OpenStudy (anonymous):

But I guessed...like I understand when I checked I did 666.67 x .15= 100 But how does 100/.15 make sense?

OpenStudy (anonymous):

Ah, alright. 75/500 =.15. This means that 15% of the people polled so far have like oranges. Assuming the 15% holds, we need an amount for which 15% of it is 100. This gives us the equation: 0.15*X = 100 Solving for X: X= 100/0.15

OpenStudy (anonymous):

I still don't get .15x100 though because I see it like out of 100 people .15 of them like oranges (the number i get) but how is it possible that its more than 100?

OpenStudy (anonymous):

Alright, so we know that 15% of people like oranges. We want to find 100 people who like oranges. So, we need to find a number such that 15% of it is 100: 15% of X = 100 0.15*X = 100

OpenStudy (anonymous):

Ok so that goes back to the finding x formula

OpenStudy (anonymous):

Great- thanks again for your help....you have made things a lot clearer to me! I appreciate it so much.

OpenStudy (anonymous):

Not a problem. Good luck! And remember, math isn't difficult, so don't make it more difficult than it needs to be!

OpenStudy (anonymous):

Ya I sometimes have an idea of how to get the answers...but I get so caught up on using the correct formula. That's why my answers are always incomplete.

OpenStudy (anonymous):

Trust yourself, and when you reach the end just ask yourself if the answer makes sense. If it does, check the work, and then move on. From what I've seen so far, you know what you're doing.

OpenStudy (anonymous):

Great- thanks for the advice. We're really lucky to have people like you to clarify things!